OpenStudy (anonymous):

Find an equation of the line containing the given pair of points: (2,0) and (-4,-3)

OpenStudy (anonymous):

formula: y-y1=slope(x-x1) slope = (y2-y1)/(x2-x1)

OpenStudy (phi):

when they give you 2 points, you can figure out m and b. (2 points define a line) slope is "change in y" divided by "change in x" take a look at this video if this doesn't ring a bell http://www.khanacademy.org/math/trigonometry/graphs/slope_of_line/v/slope-of-a-line

OpenStudy (anonymous):

ok so i have a little bit of it worked out with everything i have previously done...